Find the multiplicative inverse of 82/16 to the power of -3/4

Grade:8

1 Answers

Monoxdifly
3 years ago
If it’s 82, I’m sorry I can’t help you. But if it’s 81 like your other question, the answer is this:
(16/81)^(-3/4) = (2^4/3^4)^(-3/4) = ((2/3)^4)^(-3/4) = (2/3)^(-3) = (3/2)^3 = 3^3/2^3 = 27/8